# DoubleAgents: Interactive Simulations for Alignment in Agentic AI

### v2

@@ 6. Deployment Studies
6.1. Goal
As the in-lab study demonstrated, the simulation was helpful, and participants felt they were getting more comfortable use agentic AI help planning and executing human coordination tasks. However, would real-world seminar organizers find DoubleAgents useful and relevant? If so, how would they use it-would they plan, draft, and send invitation, follow-up, or clarification emails in different ways? Would alignment be established in a similar manner? What additional challenges might arise in real-world use? To explore these questions and evaluate the real-world deployability of DoubleAgents, we invited organizers who were managing actual events across three real-world deployments.
6.2. Deployment Participants and Procedures
@@ 6.2.1. Use Cases and Participants
In our deployment study, DoubleAgents was used in three real-world event-planning workflows involving two organizers (P11, P12). P11 used DoubleAgents to coordinate a Fall 2025 HCI seminar series (Aug 2025; Deployment 1, D1), inviting 10 speakers from a pool of 23 to fill 10 slots; we observed and logged system use for 7 days during this process. P11 later used DoubleAgents again for a Spring 2026 HCI seminar series (Jan 2026; Deployment 3, D3), inviting 11 speakers from a pool of 20 to fill 11 slots; we tracked use for 11 days. P12 used DoubleAgents to organize a conference workshop panel (Sep 2025; Deployment 2, D2), inviting 5–6 panelists from a pool of 18 to fill a single slot; we tracked use for 7 days. We recruited P11 and P12, two experienced seminar organizers (one male, one female; average age 31), via personal connections. Both were PhD students who had previously organized departmental research seminars and had prior experience using generative AI for writing emails.
6.2.2. Introduction, Setup, and Interactive Simulation Rounds
Organizers were first informed about the study and provided consent. They were then introduced to the study procedures and the DoubleAgents system through a demonstration, which included planning–execution loops, illustrated with a simulation example from the user study. Organizers also completed the initial survey covering AI perception and demographics.
To set up the system, we disabled the respondent module from the lab version of DoubleAgents and connected the SendEmail() and checkEmailUpdates() tools to Gmail’s SMTP service, enabling organizers to send emails and receive replies through their personal Gmail accounts.
For the simulation rounds, organizers prepared a simulation spreadsheet using their real organizing data (including real potential speakers’ names), first with placeholder email addresses and later with the real ones, and provided their Gmail SMTP credentials. Clear instructions and consent procedures ensured again that participants understood data usage, privacy rights, and the option to withdraw. Before connecting their real email accounts, participants completed a 2-hour interactive simulation session using pre-generated speaker personas based on the real candidates, allowing them to safely familiarize and themselves align with the system’s workflows. Finally, they tested the real email-sending module twice by sending messages to placeholder addresses (the researchers’ and organizers’ personal emails) to confirm successful delivery; afterward, we switched the recipient addresses to the invitees’ real Gmail accounts.
6.2.3. Live Email Execution and Response Handling
After the simulations, organizers used DoubleAgents to initiate the coordination process and track it for over a few days. They continued receiving system updates over the following week (D1: 7 days, D2: 4 days, D3: 11 days). We communicated with participants consistently, logging interactions, usage patterns, and feedback on the system’s support for scheduling, tracking, and learning. Organizers’ system interactions, including email sends, scheduling updates, and system adaptations, were tracked. Qualitative feedback was collected via check-in meetings, while quantitative system logs were collected for system learning over the week.
@@ 6.3. Deployment Results
6.3.1. DoubleAgents’ Perceived Usefulness and Relevance in Real-Life Human Coordination Problems.
After the interactive simulations, both organizers (P11 and P12) reported that DoubleAgents was highly relevant to their existing seminar-organizing workflows. They highlighted that the system fit seamlessly into their email-heavy tasks, and that its control settings and edge-case flagging helped them handle complex coordination scenarios that conventional scheduling tools (e.g., Calendly) cannot easily manage. P11 emphasized that crafting good email templates is a delicate and important task because it requires personalizing messages to “make the speakers feel they are special,” even when requesting a favor. Both P11 and P12 found DoubleAgents capable of generating strong starter emails for them to edit early on, as well as producing acceptable final emails to send. P12 especially appreciated having explicit coordination policies in the system to guide next-step decisions in ways that matched his decision-making habits and values. He also found it useful that rationales were always visible and editable, in contrast to AI tools that hide their reasoning. After using DoubleAgents, P12 remarked, “I trust AI [systems] for specific use cases more,” explaining that he found DoubleAgents dedicated to and capable of generating high-quality emails. P11 found the color-coded visualization of speaker constraints and response states (e.g., not yet outreached, tentative yes, yes, no) especially useful; after D1, she adopted a similar color-coded scheme for other tracking tasks she previously managed in spreadsheets.
6.3.2. Interactive Simulation as an Alignment Method for Complex, Realistic Tasks.
Both P11 and P12 emphasized that interactive simulation was important for understanding DoubleAgents and for setting up initial coordination policies and email templates. They appreciated that simulations used real invitee names and incorporated persona cues and edge cases (e.g., funding requests, cancellations, Zoom requests), which helped them anticipate challenges and craft templates that would transfer to live deployment. For example, P12 used early simulation rounds to iterate on what information to include and where to place key details (e.g., the workshop webpage link) in his email templates, similar to how he had previously tested invitation templates by emailing himself.
P11’s D1 simulation surfaced multiple policy- and tone-sensitive situations (e.g., rescheduling after a slot had already been assigned), which prompted iterative template and policy revisions. P11 reported that the system’s drafts were largely aligned with her language, but she removed or adjusted certain lines to better match her preferences and to clarify what information should be included when edge cases occur. Overall, P11’s D1 interactive simulation took four days across five meetings: the first two meetings focused on ensuring the system ran smoothly for a high-complexity scenario (20+ candidates, 10 slots, and substantial context per speaker). During setup, P11 revised invitation and follow-up templates more than four times and updated speaker descriptions so email generation could incorporate richer context. After we refined the system and onboarding protocols, P12’s D2 simulation took two days across two meetings. P12’s panel invitation task was also simpler (e.g., shorter descriptions and a single-slot panel), which further reduced the needed time. Both organizers reported that practice simulation rounds were valuable, and that sending test emails via SMTP to his own personal gmail helped him gradually understand system capabilities and limitations. Overall, these deployments suggest that simulation and onboarding effort scales with task complexity and the amount of required context.
During D1, P11 specifically highlighted the importance and difficulty of providing sufficient context for highly personalized emails, especially for real-world cases. She noted that “[persona description management] is really tricky.” For instance, one invitee had previously coordinated with a different organizer in a prior semester (Friday availability), but the D1 seminar series had shifted to Thursdays. To avoid missing or conflicting context across past threads, schedule changes, and templates, P11 updated the persona description used for simulation with an additional line noting that the invitee had been told the wrong day in the past (Friday 4–5pm) and that the seminar was now on Thursday 4–5pm. In another case, P11 had already discussed the seminar with a speaker on Discord but initially omitted this context during early simulations; she then added it to the persona description so the follow-up email could adopt the appropriate tone and content. After rehearsing with real-speaker data in simulation, P11 finalized a preferred template for live sending.
P11 also noted that even in D3 (her second deployment), simulation remained necessary because “real life is super complex.” She observed new patterns in edge cases, such as delayed responses or vague “check back next month” replies from senior faculty, which led her to revise policies (e.g., longer follow-up intervals) and adopt a more formal email tone. Across both deployments, simulation served not just as a testing ground, but as a means of progressive alignment: helping users adapt policies, templates, and expectations to shifting social and contextual dynamics. These findings suggest that onboarding and setup effort scale with task complexity and that simulation continues to play an ongoing role in supporting user-aligned adaptation, even after initial deployment.
6.3.3. Example: Live Email Execution, Response Handling, and Edge Cases
During D3, P11 sent emails to 11 selected speakers and relied more heavily on templates and system-generated drafts on Day 1 (a Friday), reflecting increased familiarity with the workflow. DoubleAgents then suggested waiting for responses over the following two days due to the weekend, which P11 accepted. Late on Day 3 (Sun), Invitee J and Invitee A replied with questions, which DoubleAgents surfaced in P11’s Day 4 (Mon) DoubleAgents session. On the morning of Day 4, Invitee B replied with availability. Based on these responses, DoubleAgents suggested replying first to Invitees A and J, but not to Invitee B, to preserve future scheduling flexibility and avoid potential conflicts. P11 accepted this suggestion and drafted replies accordingly. Specifically, Invitee A asked a funding-related question and included a personal greeting (e.g., “Happy New Year… if it doesn’t materialize…”). P11 revised the AI-generated reply by incorporating similar phrasing from their own writing style, but chose not to save this edit as a reusable template, noting that it reflected a personal, context-specific language choice. On Day 5, Invitees I and J replied, and DoubleAgents suggested allocating them to earlier-semester slots, which P11 accepted and responded to. Similar reply-and-assign cycles occurred on Day 6. On Day 7, DoubleAgents suggested following up with all the invitees who had not yet responded, in accordance with P11’s newly updated D3 policy to follow up every four working days (rather than every two, as in D1). P11 proceeded using a personalized follow-up template that incorporated questions added on Day 4. Also on Day 7, P11 handled additional questions and confirmed slots with Invitees C and H, who had replied promptly with apologies and availability. Later that evening, Invitee I emailed to report a schedule change and requested reassignment to an unclaimed date. On Day 8, DoubleAgents suggested reassigning Invitee I and requesting additional availability from Invitees H and B to resolve emerging conflicts. P11 accepted and carried out these actions. Over the weekend (Days 9 and 10), Invitee H responded with new dates but noted that P11 should follow up again the following month, as their schedule was still tentative. DoubleAgents accordingly scheduled a follow-up for Day 35, while on Day 11 it prompted P11 to address remaining unresponsive and unresolved cases.
Collectively, these interactions illustrate the complexity and multi-threaded nature of this real-world task. As P11 reflected, “Even though I do this [seminar scheduling] all the time, [DoubleAgents] was really helpful and saved me a lot of time… Policy-abiding [actions and template-following] emails are really useful. The new policies and templates I developed through this semester’s simulation—since everyone is a professor—worked really well. The simulation part cannot be skipped.” This case highlights DoubleAgents’ strong use case, demonstrating how simulation, policies, and email templates work together to support alignment between the system’s actions and users’ intentions.
@@ 7. Discussion
7.1. Human-AI Alignment via Interactive Simulations
As AI systems evolve from passive tools to agentic collaborators that plan and act on users’ behalf, human–AI alignment becomes not just a technical challenge but an interactive one. Our findings suggest that interactive simulation offers a compelling design mechanism to support this alignment by foregrounding user agency, surfacing system behavior early, and fostering a co-adaptive relationship between humans and AI.
Across our lab study and deployments, we observed that interactive simulation enabled users to engage with alignment as a gradual process, rather than a one-time configuration step. Participants entered the system unsure of their own preferences and reluctant to delegate; yet through repeated simulated cycles, each revealing system actions, uncertainties, and edge cases, they clarified values, authored and refined policies, and ultimately built enough user comfort to let the system act on their behalf. This shift mirrors real-world dynamics of human delegation: people do not fully trust a new assistant immediately but gain confidence through concrete interactions over time.
Simulation thus serves multiple alignment functions simultaneously. First, it externalizes system behavior, making invisible decision logic (e.g., why the agent followed up or chose a certain speaker) legible to users. Second, it creates safe interaction spaces, akin to usability testing for AI plans. Like sandboxed environments in traditional software testing, simulations help users anticipate downstream consequences without real-world stakes, crucial in settings like email coordination where tone, timing, and etiquette are socially sensitive. Third, simulation supports user-side alignment, helping users identify preferences they hadn’t articulated upfront (e.g., when to prioritize seniority, how long to wait before following up, or what tone feels appropriate across contexts). Rather than requiring exhaustive foresight, DoubleAgents invites users to iterate on policies as new situations emerge.
This design reframes alignment as a human-AI collaboration process: the agent proposes actions, but the human remains in the loop, interpreting, approving, and updating the future behavior of the system. In contrast to approaches that aim to extract user goals from demonstrations or feedback, we emphasize the value of direct, interactive co-construction of alignment artifacts, such as stop hooks, policy rules, and email templates, which evolve with use cases. This opens new design directions: how can we better scaffold early-stage alignment for users who don’t yet know their own preferences? What kinds of interfaces help users develop trust while retaining oversight? How might simulation be extended to other high-stakes or interpersonal agentic domains, such as hiring, negotiation, or caregiving? Our work suggests that alignment is not a prerequisite to agentic AI, but a continuous, situated interaction practice. By embedding simulation into the human–AI workflow, we give users the tools not just to observe and react to agent behavior, but to actively shape and steer it. In doing so, DoubleAgents contributes to a broader vision of agentic AI that is transparent, participatory, and grounded in human values from the start.
7.2. Designing User Onboarding Experiences for Agentic AI
Onboarding is a critical site for agentic AI systems. Unlike traditional tools, agentic systems require users to make early decisions about goals, decision boundaries, tone and formality, and what to delegate. They also create distinct discoverability challenges: users often do not know what the agents can do and what they can do. Key modes and safeguards (e.g., auto-send vs. draft-only) may be implicit, and failures can appear only in rare or edge cases—so there is no single, linear “progress bar” that reveals system status or readiness. As a result, effective onboarding is necessary for users to form accurate mental models, judge reliability, and feel comfortable delegating and deploying the agent in higher-stakes settings (e.g., a seminar organizer connecting to Gmail to send real invitations to senior faculty—who may also be potential postdoc advisors—where mistakes can harm professional relationships and reputation). In DoubleAgents, we found that interactive simulation provides a powerful onboarding scaffold, helping users observe, evaluate, and refine system behavior before real-world use.
Rather than presenting configuration as a static setup phase, DoubleAgents embeds onboarding through repeated rounds of interactive simulation. During these rounds, edge cases (e.g., a speaker canceling or requesting Zoom) expose users to socially complex scenarios, help them become familiar with the system’s behavior, and elicit their personal preferences. Participants in our study often hesitated to delegate initially, not due to lack of interest, but due to uncertainty about agent behavior. Using repeated rounds of interactive simulation as an onboarding task lowered this barrier by helping participants move past the initial learning curve and novelty effects, and by revealing how the agent behaved across diverse cases. Like training a new assistant, onboarding becomes a co-configuration process in which expectations, autonomy, perceived usefulness, and trust are developed together.
We argue that onboarding for agentic AI should not be treated as a brief tutorial animation or tooltip overlay. Instead, it can be designed as a structured, interactive process where users and systems collaboratively define goals, boundaries, and escalation rules. DoubleAgents shows how early simulations can accelerate alignment, ease delegation, and set the foundation for long-term, trustworthy collaboration.
@@ 8. Limitations and Future Work
Our work has several limitations that point to promising directions for future research. First, DoubleAgents was initially evaluated through simulated seminar-organizing scenarios. While these simulations were reviewed by domain experts and designed to reflect realistic coordination challenges, they cannot fully capture the diversity and unpredictability of real-world interactions. Future work could strengthen alignment by grounding both the coordination agent and respondent simulations in users’ real historical data, such as past email threads, organizing decisions, and communication patterns. This would enable more personalized modeling of users’ values, tone, and escalation preferences, as well as more behaviorally faithful respondent agents. Second, DoubleAgents currently relies on different language models for different subtasks (e.g., o3 for reasoning and 4o-mini for generation). Although our technical evaluation suggests this division performs adequately, future models with stronger social reasoning and long-horizon planning capabilities may further improve coordination quality. In addition, tighter integration between user-specific data and model reasoning, such as learning when and how to apply policies from prior interactions, remains an open challenge. Third, while we conducted real-world deployments with two seminar organizers, discussions with additional organizers revealed coordination scenarios that our current system does not fully support. For example, bilingual invitations, multi-organizer decision-making, and negotiation among organizers introduce shared authority, language switching, and conflicting preferences. Extending DoubleAgents to support collaborative alignment, where policies, stop hooks, and responsibilities are jointly defined and negotiated, represents an important direction for future work. Finally, all study participants were experienced users of AI tools, which may have positively influenced their willingness to delegate and their trust in the system. Future studies should include participants with a wider range of AI familiarity to better understand onboarding challenges, mental model formation, and how interactive simulation can support alignment for users with less technical background.

### v3

@@ 6. Deployment Studies
6.1. Study Design
6.1.1. Goal
The lab study demonstrated that interactive simulation helped participants become more comfortable delegating socially embedded tasks to agentic AI. The deployment study examines whether this alignment enables organizers to confidently offload real coordination tasks, and how alignment evolves when the agent operates in live, socially complex environments.
@@ 6.1.2. Use cases and participants
DoubleAgents was used in three real-world event-planning deployments (D1–D3) involving two organizers (Q1, Q2), both PhD students (one male, one female; average age 31) with prior experience organizing departmental seminars and using generative AI for email writing. Q1 coordinated two HCI seminar series (D1, D3) and Q2 organized a conference workshop panel (D2). Table 2 summarizes each deployment.
Deployment ID D1 D2 D3
Organizer ID Q1 Q2 Q1
Task Fall seminar Workshop panel Spring seminar
Date Aug 2025 Sep 2025 Jan–Feb 2026
Candidate pool 23 18 20
Speakers invited 10 6 11
Deployment length 7 days 4 days 12 days
Table 2. Summary of three real-world deployments.
6.1.3. Deployment procedure
Organizers were first informed about the study and provided consent, then introduced to the system through a demonstration of planning–execution loops using a simulation example from the lab study. Before going live, each organizer completed at least two interactive simulation sessions (2+ hours) using personas based on their real candidates, iterating on policies, templates, and contextual inputs until the agent’s proposed actions matched the organizer’s intent. For deployment, we disabled the respondent module and connected the system to Gmail’s SMTP service, enabling organizers to send and receive real emails through their personal accounts. Once organizers judged their alignment sufficient, they used DoubleAgents to initiate and manage their coordination tasks over the live durations listed in Table 2. During execution, the agent operated on live emails while organizers retained approval authority over each action. We tracked system interactions including email sends, scheduling updates, and policy adaptations, and collected qualitative feedback via regular check-in meetings.
@@ 6.2. Deployment Results
6.2.1. Alignment as a precondition for offloading.
Both organizers reported that DoubleAgents fit seamlessly into email-heavy coordination workflows that conventional tools (e.g., Calendly) cannot manage, and attributed their willingness to offload tasks to the system’s alignment mechanisms—explicit, editable policies and templates that made the agent’s reasoning legible. Q2 valued having editable policies to guide next-step decisions, contrasting this with AI tools that hide their reasoning; after deployment, Q2 remarked, “I trust AI [systems] for specific use cases more.” Q1 highlighted that messages must “make speakers feel they are special” and found DoubleAgents capable of generating strong starter drafts that improved over time.
6.2.2. Building alignment through and beyond simulation.
Both organizers used interactive simulation with real invitee names and persona-specific edge cases (e.g., funding requests, cancellations, Zoom preferences) to iteratively refine policies and templates before going live. Q2 used early rounds to iterate on template structure; Q1’s D1 surfaced tone-sensitive situations (e.g., rescheduling after a slot was already assigned), prompting revision of both templates and policies. Setup effort scaled with task complexity: Q1’s D3 (20 candidates, 11 slots, substantial per-speaker context) required four days across five meetings; Q2’s simpler panel invitation took two days across two meetings. A recurring challenge was context curation—Q1 noted that persona description management was “really tricky,” requiring updates such as clarifying schedule changes or adding context from prior Discord conversations so emails would adopt the right tone. Simulation-established alignment transferred effectively to structurally predictable scenarios—funding questions, scheduling conflicts, and reassignment requests were handled with only minor adjustments—but deployment surfaced unanticipated dynamics: noncommittal faculty responses (e.g., “check back in two months”) and cross-platform context required ongoing policy and tone adjustments. Even in her second deployment, Q1 found renewed simulation necessary because “real life is super complex.
6.2.3. Distributed cognition across organizer and agent.
Beyond alignment, the deployments revealed how DoubleAgents redistributed cognitive labor between organizer and agent through three complementary mechanisms:
Visualization as shared group representation.
Over deployment, Q1 found DoubleAgents’ color-coded visualization of speaker states (e.g., tentative yes, confirmed, declined) so useful that she adopted a similar color-coded scheme for other social tracking tasks she managed in spreadsheets beyond D1 and D3. Q2 drew on his existing practice of using ranked spreadsheets to monitor invitation progress, and described DoubleAgents’ visualization as supporting trust through “natural mapping” to that workflow.
Policy and artifact refinement as externalized strategy.
Over deployment, both organizers iteratively edited policies and templates to improve alignment. Q1’s D1 simulation surfaced policy-sensitive situations (e.g., rescheduling after a slot had already been assigned), prompting new policies. For D3—targeting more senior faculty than D1—she changed follow-up frequency from every two working days to every four, adopted more formal language, and revised templates more than four times to match the tone she identified after D1. Q2 valued explicit policies that matched his decision-making habits, appreciated that rationales were always visible and editable (unlike AI tools that hide reasoning), and drafted new policies to address gaps found during simulation. During deployment, Q1 observed that her edits carried forward to subsequent emails, so the artifacts accumulated organizational knowledge over time, requiring fewer changes. After one initial revision, Q2 showed a similar pattern: over usage, he began skipping full review of longer agent-drafted emails, trusting the style transfer and its consistency.
Coordination agent as attention manager.
In Q1’s D1 and D3 deployments, DoubleAgents managed parallel reply-and-assign cycles across 11 speakers over two weeks, at times handling three or four near-simultaneous replies with overlapping availability conflicts. Even in Q2’s smaller D2, he found it sufficient to check DoubleAgents for coordination status rather than scanning his inbox. Q1 shared that DoubleAgents saved her time at the end of D3: “Even though I do this [seminar scheduling] all the time, [DoubleAgents] was really helpful and saved me a lot of time… Policy-abiding [actions and template-following] emails are really useful. The new policies and templates I developed through this semester’s simulation—since everyone is a professor—worked really well.
@@ 7. Discussion
7.1. Beyond Organizing Seminars
While DoubleAgents is instantiated for seminar organization, its underlying architecture is domain-general. Extending it to other coordination settings highlights new design opportunities across each component of the distributed cognition framework.
Bespoke visualizations for different domains. Our dashboard provided temporal, social, and procedural views tailored to scheduling. Other domains will require different external representations: a hiring pipeline may benefit from a funnel view across candidate stages, while conference planning may require dependency graphs linking venues, catering, and speaker logistics. The insight from distributed cognition is that the right external representation enables evaluation at a glance (Hutchins and Klausen, 1996), which means the visualization layer should be treated as a bespoke design problem for each domain. One promising direction is to make this layer adaptive: users could specify what they need to monitor, and the system could generate or iteratively refine visual layouts. Such a malleable representation layer would allow the interface to co-evolve with users’ understanding of the task.
From single-step policies to multi-step processes. Our current policy module encodes single-step conditionals (e.g., when X happens, do Y). However, many coordination tasks unfold as multi-step procedures. For example, inviting a senior speaker may involve an informal inquiry, followed by a formal invitation, and then logistical coordination, each governed by distinct norms. Supporting such process-level specifications would align the system more closely with how organizations encode knowledge: as structured workflows or playbooks rather than isolated rules (Hutchins and Klausen, 1996). This shift also introduces new interface challenges: users must be able to track not only which policy is active, but where each stakeholder sits within an evolving process.
Integration with existing coordination ecosystems. Real-world coordination is deeply embedded in existing tools: communication occurs in Slack, tasks in Jira, schedules in shared calendars, and budgets in spreadsheets. Integrating DoubleAgents with these systems (e.g., via APIs or MCP servers) would allow the agent to both read from and act within the user’s existing infrastructure. For instance, the agent could summarize threads to Slack, create follow-up tasks, or detect stalled interactions from calendar and messaging signals. Beyond convenience, such integration enriches the agent’s state representation, enabling it to reason over richer, continuously updated signals and reducing the burden on users to manually maintain context.
7.2. Generalizing to Multi-User Coordination
Our study focuses on a single organizer, but many real-world coordination tasks involve multiple stakeholders with distinct roles. Distributed cognition naturally extends to such settings, where cognition is shared across people and artifacts rather than confined to an individual (Hutchins and Klausen, 1996). This raises new design considerations for how DoubleAgents might support team-based coordination.
Role-specific views with shared context. In multi-user settings, each participant requires a role-specific interface while remaining aware of cross-role dependencies. For example, in workshop organization, one person may manage catering, another speaker invitations, and another the budget. Each role demands a tailored view (e.g., headcounts, speaker pipeline, financial tracking), yet these views are tightly coupled: budget constraints venue choice, and venue capacity constrains invitations. A multi-user DoubleAgents system would provide each user with their own dashboard, policies, and agent, while maintaining a shared coordination substrate that propagates updates across roles. Changes in one domain (e.g., new funding or speaker cancellations) would be surfaced as actionable signals in others, enabling coordinated adaptation without requiring manual information relay.
Per-user and shared policies. Delegation preferences vary across individuals. One user may permit autonomous actions within their domain, while another may require explicit approval for every step. Supporting this heterogeneity requires both per-user policies, which govern behavior within individual scopes, and shared policies, which encode team-level constraints (e.g., approval requirements for large expenses). In this setting, stop hooks evolve into coordination mechanisms not only between a user and their agent, but across users: they can flag when an action in one role impacts another and trigger cross-role review before execution. This enables alignment to be negotiated at both individual and collective levels.
From individual to team alignment. Our findings suggest that alignment improves through repeated interaction. In multi-user settings, this process becomes collective: teams develop shared expectations about when the system acts autonomously, when it defers, and how it communicates state. Over time, accumulated policies, templates, and stop hooks form a persistent layer of organizational memory: an encoded representation of how the team coordinates (Hutchins and Klausen, 1996). Similar to how teams develop shared mental models through repeated use of tools and procedures, iterative interaction with such systems can produce reusable coordination practices that transfer across team members and tasks.
@@ 8. Limitations and Future Work
Our lab study involved 10 participants in simulated scenarios, limiting generalizability, though the within-subjects design across two sessions yielded rich insights into how alignment strategies and delegation boundaries evolved with repeated use. Three week-long deployments validated these patterns in real-world settings. Larger, more diverse samples, including users with less AI familiarity, would help clarify onboarding challenges and how simulation supports alignment across experience levels. While our simulations reflected realistic coordination challenges, they cannot fully capture real-world unpredictability. Grounding both the coordination agent and respondent simulator in users’ historical data, including past emails, decisions, and communication patterns, could enable more personalized modeling of values and tone. Finally, DoubleAgents relies on different language models for different subtasks (e.g., o3 for reasoning, 4o-mini for generation). Future models with stronger social reasoning and long-horizon planning may improve coordination quality, and tighter integration between user-specific data and model reasoning remains an open challenge.